# Region Guided Attention Network for Retinal Vessel Segmentation

### v1

@@ Abstract
Retinal imaging has emerged as a promising method of addressing this challenge, taking advantage of the unique structure of the retina. The retina is an embryonic extension of the central nervous system, providing a direct in vivo window into neurological health. Recent studies have shown that specific structural changes in retinal vessels can not only serve as early indicators of various diseases but also help to understand disease progression. In this work, we present a lightweight retinal vessel segmentation network based on the encoder-decoder mechanism with region-guided attention. We introduce inverse addition attention blocks with region guided attention to focus on the foreground regions and improve the segmentation of regions of interest. To further boost the model’s performance on retinal vessel segmentation, we employ a weighted dice loss. This choice is particularly effective in addressing the class imbalance issues frequently encountered in retinal vessel segmentation tasks. Dice loss penalises false positives and false negatives equally, encouraging the model to generate more accurate segmentation with improved object boundary delineation and reduced fragmentation. Extensive experiments on a benchmark dataset show better performance (0.8285, 0.8098, 0.9677, and 0.8166 recall, precision, accuracy and F1 score respectively) compared to state-of-the-art methods.
@@ 1 Introduction
As a direct extension of the nervous system, the eye is the only part of the body where the microneuronal and microvascular system can be viewed noninvasively, providing an accessible method for diagnosing and monitoring the effects of systemic diseases and drugs. This provides clinicians with a potential method for using the eyes to diagnose and monitor systemic disease and to assess the impact of systemic disease on eye health. Recent clinical investigations have highlighted the potential of using retinal imaging biomarkers to detect not only eye diseases (such as glaucoma and age-related macular degeneration), but also various other diseases, including hypertension, dementia, Parkinson’s disease and multiple sclerosis, in their preclinical and presymptomatic stages. In addition, it also helps diagnose conditions related to brain and heart health, which exhibit abnormal variations in the vascular structure of the retina. Therefore, accurate segmentation of the retinal vessels makes it possible to build an automated diagnosis system, and the segmentation of the retinal vessels has attracted the interest of researchers.
@@
Accurate segmentation of retinal vessels is hampered by the challenges posed by image features such as low contrast and imbalanced intensity, and anatomic features such as variations in thickness of the main vessels and capillaries. In addition, the presence of exudates and lesions in the retinal fundus image further complicates the segmentation task. To overcome these hurdles, researchers have employed a range of supervised or unsupervised algorithms alongside computer vision techniques, aiming at accurate and automated segmentation [1, 2]. Recent advances indicate that deep learning architectures outperform other methodologies in this domain [3, 4, 5]. Therefore, various deep learning strategies have contributed to the advancement of retinal vessel segmentation.
@@
U-Net [6], originally designed for medical image segmentation, exhibits a drawback in the identification of false boundaries within retinal images alongside blood vessels. Yan et al. [7] bolstered U-Net’s efficacy by implementing segment-level loss which emphasises the thickness consistency of thin vessels. Gu et al. [8] proposed a context encoder to capture high-level features and used pre-trained ResNet blocks to improve retinal vessel segmentation. Wang et al. [9] introduced DEU-Net, which employs a fusion module function to merge a spatial path with a large kernel. This integration preserves spatial data while capturing semantic details effectively. Dulau et al. [10] developed a postprocessing pipeline named VNR (Vessels Network Retrieval) to ensure a connected structure for retinal vessel networks, improving segmentation accuracy by removing misclassified pixels and reconnecting disconnected branches. Fu et al. [11] used a multiscale, multilevel convolutional neural network (CNN) to obtain a dense hierarchical representation and also incorporated a conditional random field (CRF) to model extended interactions among pixels. However, despite the efficacy of these methods, they overlook the necessity to optimise computational efficiency to adapt the network for use in resource-limited embedded systems.
@@
Recently, researchers have shown increased interest in lightweight networks for the segmentation of general objects as well as medical images [12, 13, 14, 15, 16]. SegNAS3D [17] introduced a framework that searches for automated network architectures for 3D image segmentation, utilising a learnable directed acyclic graph representation to optimise hyperparameters and achieve superior segmentation results with reduced computational cost and smaller network architectures compared to manual approaches. IC-Net [18] introduced an image cascade network that effectively reduces computation for real-time semantic segmentation and accelerates model convergence. Xception [19] and MobileNet [20] use depth-wise separable convolutions to reduce the parameter count and computational complexity, making them suitable for devices with limited computational resources. They both improve performance on and efficiency in image classification and segmentation.
@@
    We present a lightweight region guided segmentation network with only 40K parameters that can be deployed on devices with limited computational and memory resources.
@@
    We introduce region-guided inverse addition attention blocks alongside weighted dice loss particularly crafted for retinal vessels segmentation that explicitly focus on foreground regions, resulting in better segmentation of the ROI.
@@
    To refine initial segmentation maps to obtain refined segmentation, we propose a partial decoder and use multiple attention blocks that align the high and low-level features.
@@
    We have performed extensive experiments and identified the best hyperparameters for segmentation of retinal vessels on benchmark datasets.
@@ 2 Related Work
Retinal vessel segmentation is a crucial task in ophthalmic image analysis, enabling the detection and monitoring of various eye diseases such as diabetic retinopathy, glaucoma and hypertensive retinopathy. Over the years, numerous approaches have been proposed that use both traditional image processing techniques and advanced machine learning algorithms. Earlier methods for segmentation of the retinal vessels that relied primarily on traditional image processing typically involve several steps, including pre-processing, vessel enhancement, segmentation and post-processing. To improve segmentation performance, some researchers used preprocessing techniques to enhance the image quality. However, advanced machine learning-based techniques can automatically learn features from large datasets, often outperforming traditional approaches.
@@
CNN based techniques have exhibited promising performance on segmentation tasks and therefore gained popularity. Uysal et al.[21] proposed a CNN architecture in an end-to-end learning framework for medical image segmentation. They used data augmentation to enlarge the dataset synthetically for better performance, yet their work faces the challenge of dependency on the dataset and the model overfits on a majority of medical image datasets due to large model capacity and small dataset size. Oliveira et al.[22] used a fully convolutional neural network (FCN) for the task of segmenting retinal vessels from fundus images. FCNs are particularly well suited for this task because of their ability to produce pixel-wise predictions, which is crucial for accurately delineating the thin and intricate structures of retinal vessels. Yan et al.[23], proposed a three stage FCN architecture to progressively refine predictions through multiple stages, each stage building on the output of the previous one. In the first stage the model gives a coarse prediction, in the second stage the coarse predictions are refined by focussing on medium-level features and improving the resolution, while in the final stage, the model achieves high precision in segmentation by correcting small errors and adding more fine detail. The three-stage FCN network proved to be very effective for the delineation and segmentation of retinal vessels. Guo et al. proposed the BTS-DSN model [24], which incorporates auxiliary supervision signals at multiple intermediate layers. This approach aids in facilitating gradient flow during the training process, leading to more stable convergence and improved overall segmentation performance. The network architecture includes short connections, akin to those found in residual networks. These connections are crucial for effectively propagating information across different layers, thus improving feature extraction and improving segmentation accuracy. The authors use ResNet-101 as the backbone of the BTS-DSN model, providing it with substantial capacity. However, this choice also renders the model computationally intensive and resource-demanding, which may be a consideration for practical deployment. Arsalan et al.[25] proposed an AI-based semantic segmentation architecture tailored for the analysis of retinal images, which leverages a deep learning framework with multiple CNN layers to achieve high precision in identifying and segmenting retinal regions affected by diabetic and hypertensive retinopathy.
@@
U-Net[6] achieved ground-breaking results on image segmentation tasks and since then, researchers have come up with numerous variations of the popular encoder-decoder-based architecture and have improved the segmentation accuracy on general object segmentation tasks, as well as on medical image segmentation. Oktay et al.[26] added attention gates to the standard U-Net to achieve better model sensitivity and segmentation accuracy. The authors combined the attention mechanism with the U-Net architecture for the task of segmenting multiclass medical images and obtained promising results while keeping low computational complexity. Jin et al.[27] introduced the DUNet architecture, which integrates deformable convolutional networks into the U-Net framework. The architecture was designed to capture more complex vessel structures and improve segmentation accuracy by adapting the receptive fields to the shape of the retinal vessels. Traditional convolutional layers have fixed geometric structures, which can be limiting when dealing with the irregular shapes of retinal vessels, while deformable convolutions address this by allowing the network to learn offsets for the convolutional kernels, enabling adaptive and flexible receptive fields that can better capture the variability in vessel shapes and sizes. Reza et al.[28] introduced the use of Bidirectional Convolutional Long-Short-Term Memory (Bi-ConvLSTM) layers within the U-Net architecture. ConvLSTM layers are designed to handle spatial and temporal dependencies in data, making them suitable for tasks that require contextual understanding over sequences or spatially dependent structures. By using Bi-ConvLSTM, the model can capture dependencies in both forward and backward directions, enhancing its ability to model complex spatial relationships.
Figure 1: Block diagram of the proposed methodology.
@@ 2.1 Light-weight Models for Medical Image Segmentation
After the success of light-weight models such as Mobile-Net [29] on general object segmentation, researchers have recently been interested in developing light-weight networks for the segmentation of medical images. They have tried to minimise the network size and capacity, decrease the overall number of computations performed, and reduce the memory occupied by the model. Iqbal et al.[30] proposed a lightweight, compact and efficient network called LDMRes-Net based on dual multiscale residual blocks that incorporate a multiscale feature extraction mechanism, allowing it to capture details at various levels of granularity. They reduced the number of parameters and computational complexity compared to traditional deep learning networks. The efficiency of the LDMRes-Net is enhanced by the implementation of depth wise separable convolutions while the residual connections in the network maintain its performance. Tariq et al. [31] have proposed a lightweight network for medical image segmentation. It focusses on capturing high-frequency features necessary for medical image segmentation tasks, and the implementation of expand and squeeze blocks makes their model robust and computationally efficient. The authors have attempted to provide a solution for applications on devices with limited computational resources. Li et al.[32] utilised a lightweight version of U-Net that is designed to be computationally efficient while maintaining high accuracy in the segmentation of lesions in ultrasound images. There is not much work done on the utilisation and introduction of lightweight models for segmentation of retinal vessels. In this paper, we focus on building a lightweight model for segmentation of retinal vessels while maintaining state of the art segmentation performance.
@@ 3 Region Guided Attention Network
In this work, we propose a region-guided attention network that uses the strengths of U-Net[6] and a region-guided attention mechanism for the segmentation of medical images. As a base model, we first modified the U-Net architecture to its lightweight version. To do so, we minimise the number of learnable parameters by reducing the number of layers and the number of filters in each layer. The motivation behind this step was to deal with the overfitting issue that occurs mainly due to lar ge model capacity and the small size of the medical images dataset that specifically deals with connected vessels. This choice of parameters makes the model computationally efficient while maintaining satisfactory segmentation performance and deals with weak anti-noise interference ability specifically for capillary vessels. To boost the segmentation performance even further, we introduce an Inverse Addition Attention mechanism that forces the model to focus on the ROI that are most relevant to the segmentation task. As shown in Figure 1, the initial segmentation map is generated by the partial decoder, and through multiple attention blocks, we refine the initial segmentation maps until we obtain the final refined segmentation map.
@@ 3.1 Encoder Block
The encoder block in the proposed model comprises two convolutional layers, each followed by a batch normalisation layer and a ReLU activation function, and is concluded with a non-overlapping max pooling operation. The primary goal of this block is to capture and refine relevant features before passing them on to subsequent blocks. To enhance computational efficiency and avoid redundant operations, we employ depthwise separable convolutions. This approach not only accelerates model training and inference, which is critical for real-time applications, but also significantly reduces the number of parameters, resulting in lower memory usage. Depthwise convolutions process each channel independently, allowing efficient spatial feature extraction without excessive parameter overhead. To achieve a balanced integration of spatial and channel-wise features, we combine depth-wise convolutions with point-wise ($1\times 1$) convolutions. This combination enables the model to effectively capture detailed and relevant features, leading to improved representation learning and overall model performance. In addition, the use of batch normalisation helps stabilise and accelerate the training process by normalising the activations of each layer. This reduces internal covariate shift and enables the use of higher learning rates, further speeding up convergence. The ReLU activation function introduces nonlinearity into the model, allowing it to learn complex patterns and interactions within the data. The nonoverlapping max pooling operation reduces the spatial dimensions of the feature maps, thereby decreasing the computational load while preserving essential spatial information.
Overall, the design of the encoder block, with its separable convolutions in-depth, batch normalisation, ReLU activations and max pooling, ensures efficient and effective feature extraction. This combination enhances the model’s ability to learn rich representations from the input data, contributing to improved performance in downstream tasks.

### v3

@@ Abstract
Retinal imaging has emerged as a promising method of addressing this challenge, taking advantage of the unique structure of the retina. The retina is an embryonic extension of the central nervous system, providing a direct in vivo window into neurological health. Recent studies have shown that specific structural changes in retinal vessels can not only serve as early indicators of various diseases but also help to understand the progression of the disease. In this work, we present a lightweight retinal vessel segmentation network based on the encoder-decoder mechanism with region-guided attention. We introduce inverse addition attention blocks with region-guided attention to focus on the foreground regions and improve the segmentation of regions of interest. To further boost the model’s performance on retinal vessel segmentation, we employ a weighted dice loss. This choice is particularly effective in addressing the class imbalance issues frequently encountered in retinal vessel segmentation tasks. Dice loss penalises false positives and false negatives equally, encouraging the model to generate more accurate segmentation with improved object boundary delineation and reduced fragmentation. Extensive experiments on a benchmark dataset show better performance (0.8285, 0.8098, 0.9677, and 0.8166 recall, precision, accuracy, and F1 score, respectively) compared to state-of-the-art methods.
@@ 1 Introduction
As a direct extension of the nervous system, the eye is the only part of the body where the micro-neuronal and micro-vascular systems can be viewed non-invasively, providing an accessible method for diagnosing and monitoring the effects of systemic diseases and drugs. This provides clinicians with a potential method for using the eyes to diagnose and monitor systemic disease and to assess the impact of systemic disease on eye health. Recent clinical investigations have highlighted the potential of using retinal imaging biomarkers to detect not only eye diseases (such as glaucoma and age-related macular degeneration), but also various other diseases, including hypertension, dementia, Parkinson’s disease, and multiple sclerosis, in their preclinical and presymptomatic stages [1, 2, 3]. In addition, it also helps diagnose conditions related to brain and heart health, which exhibit abnormal variations in the vascular structure of the retina [4, 5, 6]. Therefore, accurate segmentation of the retinal vessels makes it possible to build an automated diagnosis system, and segmentation of the retinal vessels has attracted the interest of researchers.
@@
Accurate segmentation of retinal vessels is hampered by the challenges posed by image features such as low contrast and imbalanced intensity, and anatomical features such as variations in thickness of the main vessels and capillaries. In addition, the presence of exudates and lesions in the image of the retinal fundus further complicates the segmentation task [7, 8, 1, 9]. To overcome these hurdles, researchers have employed a range of supervised or unsupervised algorithms alongside computer vision techniques, aiming for accurate and automated segmentation [10, 11]. Recent advances indicate that deep learning architectures outperform other methodologies in this domain [12, 13, 14]. Therefore, various deep-learning strategies have contributed to the advancement of retinal vessel segmentation.
@@
U-Net [15], originally designed for medical image segmentation, exhibits a drawback in the identification of false boundaries within retinal images alongside blood vessels. Yan et al. [16] bolstered U-Net’s efficacy by implementing segment-level loss which emphasises the thickness consistency of thin vessels. Gu et al. [17] proposed a context encoder to capture high-level features and used pre-trained ResNet blocks to improve retinal vessel segmentation. Wang et al. [18] introduced DEU-Net, which employs a fusion module function to merge a spatial path with a large kernel. This integration preserves spatial data while effectively capturing semantic details. Dulau et al. [19] developed a post-processing pipeline named VNR (Vessel Network Retrieval) to ensure a connected structure for retinal vessel networks, improving segmentation accuracy by removing misclassified pixels and reconnecting disconnected branches. Fu et al. [20] used a multiscale, multilevel convolutional neural network (CNN) to obtain a dense hierarchical representation and also incorporated a conditional random field (CRF) to model extended interactions among pixels. However, despite the efficacy of these methods, they overlook the need to optimise computational efficiency to adapt the network for use in resource-limited embedded systems.
@@
Recently, researchers have shown an increased interest in lightweight networks for the segmentation of general objects and medical images [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. SegNAS3D [31] introduced a framework that searches for automated network architectures for 3D image segmentation, utilising a learnable directed acyclic graph representation to optimise hyperparameters and achieve superior segmentation results with reduced computational cost and smaller network architectures compared to manual approaches. IC-Net [32] introduced an image cascade network that effectively reduces computation for real-time semantic segmentation and accelerates model convergence. Xception [33] and MobileNet [34] use depth-wise separable convolutions to reduce the parameter count and computational complexity, making them suitable for devices with limited computational resources. They both improve performance and efficiency in image classification and segmentation.
@@
    We present a lightweight region-guided segmentation network with only 40K parameters that can be deployed on devices with limited computational and memory resources.
@@
    We introduce region-guided inverse addition attention blocks along with weighted dice loss specifically crafted for retinal vessel segmentation that explicitly focuses on foreground regions, resulting in better segmentation of the ROI.
@@
    To refine the initial segmentation maps to obtain the refined segmentation, we propose a partial decoder and use multiple attention blocks that align the high- and low-level features.
@@
    We have performed extensive experiments and identified the best hyperparameters for the segmentation of retinal vessels on benchmark datasets.
@@ 2 Related Work
Retinal vessel segmentation is a crucial task in ophthalmic image analysis, enabling the detection and monitoring of various eye diseases such as diabetic retinopathy, glaucoma, and hypertensive retinopathy. Over the years, numerous approaches have been proposed that use both traditional image processing techniques and advanced machine learning algorithms. Earlier methods for segmentation of the retinal vessels, that relied primarily on traditional image processing, typically involved several steps including pre-processing, vessel enhancement, segmentation, and post-processing. To improve segmentation performance, some researchers used preprocessing techniques to enhance the image quality. However, advanced machine learning-based techniques can automatically learn features from large datasets, often outperforming traditional approaches.
@@
CNN-based techniques have exhibited promising performance in segmentation tasks and therefore have gained popularity. Uysal et al.[35] proposed a CNN architecture in an end-to-end learning framework for medical image segmentation. They used data augmentation to enlarge the dataset synthetically for better performance, yet their work faces the challenge of dependency on the dataset and the model overfits on a majority of medical image datasets due to large model capacity and small dataset size. Oliveira et al.[36] used a fully convolutional neural network (FCN) for the task of segmenting retinal vessels from fundus images. FCNs are particularly well suited for this task because of their ability to produce pixel-wise predictions, which is crucial to accurately delineating the thin and intricate structures of retinal vessels. Yan et al.[37], proposed a three-stage FCN architecture to progressively refine predictions through multiple stages, each stage building on the output of the previous one. In the first stage, the model gives a coarse prediction, in the second stage the coarse predictions are refined by focusing on medium-level features and improving the resolution, while in the final stage, the model achieves high precision in segmentation by correcting small errors and adding more fine details. The three-stage FCN network proved to be very effective for the delineation and segmentation of retinal vessels. Guo et al. proposed the BTS-DSN model [38], which incorporates auxiliary supervision signals at multiple intermediate layers. This approach aids in facilitating gradient flow during the training process, leading to more stable convergence and improved overall segmentation performance. The network architecture includes short connections, akin to those found in residual networks. These connections are crucial for effectively propagating information across different layers, thus improving feature extraction and improving segmentation accuracy. The authors use ResNet-101 as the backbone of the BTS-DSN model, providing it with substantial capacity. However, this choice also renders the model computationally intensive and resource-demanding, which may be a consideration for practical deployment. Arsalan et al.[39] proposed an AI-based semantic segmentation architecture tailored for the analysis of retinal images, which leverages a deep learning framework with multiple CNN layers to achieve high precision in identifying and segmenting retinal regions affected by diabetic and hypertensive retinopathy.
@@
U-Net[15] achieved ground-breaking results in image segmentation tasks and since then researchers have come up with numerous variations of the popular encoder-decoder-based architecture and have improved segmentation accuracy in general object segmentation tasks, as well as on medical image segmentation. Oktay et al.[40] added attention gates to the standard U-Net to achieve better model sensitivity and segmentation accuracy. The authors combined the attention mechanism with the U-Net architecture for the task of segmenting multiclass medical images and obtained promising results while keeping low computational complexity. Jin et al.[41] introduced the DUNet architecture, which integrates deformable convolutional networks into the U-Net framework. The architecture was designed to capture more complex vessel structures and improve segmentation accuracy by adapting the receptive fields to the shape of the retinal vessels. Traditional convolutional layers have fixed geometric structures, which can be limiting when dealing with irregular shapes of retinal vessels, while deformable convolutions address this by allowing the network to learn offsets for the convolutional kernels, enabling adaptive and flexible receptive fields that can better capture the variability in vessel shapes and sizes. Reza et al.[42] introduced the use of Bidirectional Convolutional Long-Short-Term Memory (Bi-ConvLSTM) layers within the U-Net architecture. ConvLSTM layers are designed to handle spatial and temporal dependencies in data, making them suitable for tasks that require contextual understanding over sequences or spatially dependent structures. By using Bi-ConvLSTM, the model can capture dependencies in both forward and backward directions, enhancing its ability to model complex spatial relationships. Wei et al.[43] introduced Genetic U-Net, a framework that leverages genetic algorithms for the automatic design of deep neural networks specifically tailored for the segmentation of the retinal vessels. This approach aims to optimise the network architecture without extensive manual intervention. The paper applies Neural Architecture Search (NAS) using genetic algorithms to explore and identify optimal network structures. This method systematically evolves network architectures to enhance performance, demonstrating a sophisticated use of NAS in medical imaging. Although the genetic algorithm optimises the network architecture, scaling this approach to very large datasets or real-time applications might be challenging due to the inherent computational demands and the iterative nature of the search process.
Figure 1: Block diagram of the proposed methodology. The figure depicts the overall proposed methodology, detailing the Encoder, Bottleneck, and Decoder Blocks. The detailed layers and operations of the Partial Decoder and Inverse Addition Attention blocks are given in the respective sections. The figure also explains what each symbol in the flow chart means.
@@ 2.1 Light-weight Models for Medical Image Segmentation
After the success of light-weight models such as Mobile-Net [44] on general object segmentation, researchers have recently been interested in developing light-weight networks for the segmentation of medical images. They have tried to minimise the size and capacity of the network, decrease the overall number of computations performed, and reduce the memory occupied by the model. Iqbal et al.[27] proposed a lightweight, compact, and efficient network called LDMRes-Net based on dual multiscale residual blocks that incorporate a multiscale feature extraction mechanism, allowing it to capture details at various levels of granularity. They reduced the number of parameters and computational complexity compared to traditional deep learning networks. The efficiency of the LDMRes-Net is enhanced by the implementation of depth-wise separable convolutions while the residual connections in the network maintain their performance. Tariq et al. [24] have proposed a lightweight network for medical image segmentation. It focuses on capturing high-frequency features necessary for medical image segmentation tasks, and the implementation of expand-and-squeeze blocks makes their model robust and computationally efficient. The authors have attempted to provide a solution for applications on devices with limited computational resources. Li et al.[45] utilised a lightweight version of U-Net that is designed to be computationally efficient while maintaining high accuracy in the segmentation of lesions in ultrasound images. There is not much work done on the utilisation and introduction of lightweight models for the segmentation of retinal vessels. In this paper, we focus on building a lightweight model for the segmentation of retinal vessels while maintaining state-of-the-art segmentation performance.
@@ 3 Region Guided Attention Network
In this work, we propose a region-guided attention network that uses the strengths of U-Net[15] and a region-guided attention mechanism for the segmentation of medical images. As a base model, we first modified the U-Net architecture to its lightweight version. To do so, we minimise the number of learnable parameters by reducing the number of layers and the number of filters in each layer. The motivation behind this step was to deal with the overfitting issue that occurs mainly due to the large model capacity and the small size of the medical images dataset that specifically deals with connected vessels. This choice of parameters makes the model computationally efficient while maintaining satisfactory segmentation performance and deals with weak anti-noise interference ability specifically for capillary vessels. To boost the segmentation performance even further, we introduce an Inverse Addition Attention mechanism that forces the model to focus on the ROIs that are most relevant to the segmentation task. As shown in Figure 1, the initial segmentation map is generated by the partial decoder, and through multiple attention blocks, we refine the initial segmentation maps until we obtain the final refined segmentation map.
@@ 3.1 Encoder Block
The encoder block in the proposed model comprises two convolutional layers, each followed by a batch normalisation layer and a ReLU activation function, and is concluded with a non-overlapping max pooling operation. The primary goal of this block is to capture and refine relevant features before passing them on to subsequent blocks. To enhance computational efficiency and avoid redundant operations, we employ depthwise separable convolutions. This approach not only accelerates model training and inference, which is critical for real-time applications but also significantly reduces the number of parameters, resulting in lower memory usage. Depthwise convolutions process each channel independently, allowing efficient spatial feature extraction without excessive parameter overhead. To achieve a balanced integration of spatial and channel-wise features, we combine depth-wise convolutions with point-wise ($1\times 1$) convolutions. This combination enables the model to effectively capture detailed and relevant features, leading to improved representation learning and overall model performance. In addition, the use of batch normalisation helps stabilise and accelerate the training process by normalising the activations of each layer. This reduces internal covariate shift and enables the use of higher learning rates, further speeding up convergence. The ReLU activation function introduces nonlinearity into the model, allowing it to learn complex patterns and interactions within the data. The nonoverlapping max pooling operation reduces the spatial dimensions of the feature maps, thereby decreasing the computational load while preserving essential spatial information.
Overall, the design of the encoder block, with its separable convolutions in-depth, batch normalisation, ReLU activations, and max pooling, ensures efficient and effective feature extraction. This combination enhances the model’s ability to learn rich representations from the input data, contributing to improved performance in downstream tasks.